Q=2 TC=100

Q=3 TC=120

How much is FC?


1
Expert's answer
2021-05-19T20:29:07-0400

Given,

At quantity 2 Total cost is 100

Therefore Average cost is  "\frac{100}{2}=50"

At quantity 3 Total cost is 120

Therefore Average cost is"\frac{120}{3}=40"

 Marginal cost would be"=\frac{\delta TC}{\delta Q}"

"=\frac{20}{1}=20"


Since, Total cost (TC) is the summation of the Variable cost (VC) and the Fixed cost (FC),

TC = FC + VC,    FC = TC - VC

Even at 0 quantity there will remain some Fixed cost.

Therefore FC = TC at 0 quantity

 

Since neither the variable cost nor the Total cost at 0 quantity is given therefore the problem cannot be solved.